# Local Order Average-Atom Interatomic Potentials

### v2

@@ I Introduction
The average-atom (AA) approach to modeling random alloys is based on an effective medium approximation in which the species of atoms are taken to be independent and consistent with their concentrations in a crystalline alloy. In this method, each atom is identical and typeless, with an effective atomic interatomic potential (IP) in which interactions between atoms are computed as an average over interactions that account for species identities weighted by their probability of occurrence assuming an ideal random alloy. The AA approach was originally introduced by Smith and Was [smith:was:1989], Ackland and Vitek [ackland:vitek:2009], and Najafabadi et al. [najafabadi:wang:1991], and revived by Curtin and coworkers [varvenne:luque:2016, nohring:curtin:2016]. The method has primarily been applied to alloy systems to study the effect of species concentrations on material properties [jian:xie:2020, zhao:xiong:2021, cheng:yuan:2021, seoane:farkas:2022, liang:goodelman:2023, fang:li:2024]. A recent application area for AA is the class of high entropy-alloys (HEAs) [George2019-pl] or multi-principal element alloys [Yeh2004-dy, Shittu2024-wb]. These are alloys that typically contain five or more species of approximately equal concentrations and which are of interest due to their high strength and high fracture toughness relative to traditional alloys [yeh:chen:2004, gludovatz:hohenwarter:2014, zhang:liu:2024]. Such systems require large simulation cells in order to correctly capture species fractional occupancy and hence benefit from AA, which can provide converged results with much smaller system sizes.
@@
Here we extend the AA formalism to explicitly account for LO effects by including information from partial radial distribution functions (RDFs) in the definition of the effective IP. The new local-order average-atom (LOAA) method is as efficient as AA, but correctly captures the effect of LO on alloy properties. In LOAA, the effective interaction between atoms is weighted based on the probabilities of observing pairs of species at a given distance as described by the partial RDFs, as opposed to ideal random alloy probabilities. We validate the approach by studying the elastic properties of a two-dimensional (2D) binary hexagonal crystal with Lennard-Jones (LJ) interactions, a three-dimensional (3D) Fe(1-x)/2Ni(1-x)/2Crx alloy with embedded-atom method (EAM) interactions, and martensitic phase transformations in off-stochiometric NiAl with EAM interactions. We compare the accuracy of LOAA with AA, and with “true species” (TS) computations, calculations using standard IPs with the actual atomic species identities, averaged over many realizations of the system. Overall we find LOAA to provide a more accurate estimate for the TS case than AA with no increase in computational cost except for the initial one-time generation of the partial RDFs. We observe that while the AA method can only accurately predict properties such as lattice constants, energies, and elastic constants for an ideal random alloy, the LOAA formulation correctly predicts these properties with increasing LO.
@@ II.2 Pairwise Interactions
where $V_{XY}(r)$ is the pair potential function for species $X$ and $Y$, and $r_{ij}$ is the distance between atoms $i$ and $j$.
First, we consider the random alloy case where atomic species are randomly assigned to atoms consistent with a specific set of concentrations. The average value (expectation) of the energy follows by applying the ensemble average in Eq. (3):
@@
In Eq. (7), we recover the AA formalism in which the energy of the system is expressed in terms of an effective pair potential between typeless atoms (i.e. with no specified species):
@@
In general, Eq. (9) has to be evaluated explicitly, making this approach somewhat more expensive computationally compared to the underlying pair potentials. [^3] For a system containing $n$ species, there are $n(n+1)/2$ function evaluations (accounting for symmetry). Pair potentials with a finite cut-off scale linearly with the number of atoms, $N$. Thus the AA formulation scales as $O(n^{2}N)$. However, as it is more general (and more efficient) to store pair potentials on a discrete grid that is interpolated on-the-fly at runtime, it is usually possible to pre-compute $V_{\rm eff}$ and therefore recover the same speed and scaling of the original pair potential model. The additional benefit of the AA approach, however, is that better statistics are obtained than in a calculation of a single realization of a finite-sized random alloy. One is effectively calculating the average behavior over all such possible realizations.
@@
    For a system with local order, i.e. short-range order, $\lim_{r\to\infty}\Phi_{XY}(r;\mathcal{C})=\delta(r;\mathcal{C})$, where $\delta(r;\mathcal{C})=1$ for all distances present in $\mathcal{C}$ and zero otherwise. This indicates that for atoms in $\mathcal{C}$ that are far apart, the probability that they are of species $X$ and $Y$ reverts to the independent case. [^5] We expect this limit to decay quickly with a dependence on the cutoff radius $r_{\rm cut}^{XY}$ of the pair potential $V_{XY}(r)$.
@@
where $g_{XY}(r_{ij})$ is the partial RDF between species $X$ and $Y$. In general, the RDF is defined in terms of a phase average evaluated over the relevant ensemble, [^6] We provide a general definition here for arbitrary conditions, although in practice we will only consider cases where the RDF is evaluated under static conditions at 0 K. such as the canonical ensemble [tuckerman:2010]. This phase average is evaluated for one specific species arrangement satisfying the concentration constraint in Eq. (1). It can be obtained from a molecular simulation or experiment for a given system at a given temperature $T$. Thus the left-hand side of Eq. (14) is the probability over all species arrangements for a single geometric configuration $\mathcal{C}$, whereas the right-hand side involves all geometric configurations with a single species arrangement $\mathcal{S}$. [^7] This means that each atom is assigned a species, such that Eq. (1) is satisfied, and then the phase average is over all possible arrangements of these atoms. The requirement that this equality holds is of the nature of an ergodic hypothesis. Since $\mathcal{C}$ is a system of infinite extent (see Section II.1), it is assumed that the probability over all species arrangements for a single geometric configuration snapshot at finite temperature (left-hand side of Eq. (14)) is equal to the probability over all geometric configurations for a specified set of species (right-hand side of Eq. (14)).
@@
    Computationally, the LOAA approach is not significantly slower than using the AA effective potential for the independent species case defined in Eq. (9), since $G_{XY}(R)$ can be precomputed. [^9] In the following, we will consider cases where the reference configuration is a perfect crystal. In those cases, LOAA can be made virtually as efficient as the underlying pair potential by precomputing a discrete set of $V_{\rm eff}$ functions, one for each of the discrete pair distances in the lattice up to a conservative neighbor distance allowing for atomic motion. Then during a run, the computation only requires a pair to be mapped to its reference configuration, the value of $R$ computed, and the appropriate $V_{\rm eff}$ function accessed. We note that if the only objective is to determine ground state and the partial RDFs are not known from experiments, then LOAA does not provide a computational benefit over brute force TS simulations since the RDF needs to be obtained. However once obtained, the effective LOAA potential can be used to perform a variety of computations on smaller systems efficiently.
@@ II.3 EAM-style Embedding Function
The passage for the second to third line follows since the terms $U^{X}(\rho_{i})$ and $s_{i}^{X}$ are independent. This is because the density $\rho_{i}$, given in Eq. (20), depends only on the species of the atoms surrounding atom $i$ and not on its species as indicated by $s^{X}_{i}$.
@@
In the event that the second-order term is small compared with the leading term in Eq. (24) (as for example shown for Cu-Ni alloys in [najafabadi:wang:1991]) and ${\rm Var}_{\mathcal{S}}(\rho_{i})$ is small (indicating convergence of the expansion in Eq. (22), then
@@ III Results
To explore the capabilities of the LOAA methodology, we study two example problems: 1) a 2D binary alloy modeled using an LJ potential where the degree of LO can be tuned by the choice of LJ parameters to allow for systematic comparison with the AA approach; and 2) two 3D metallic alloy systems (FeNiCr and NiAl) modeled using EAM potentials where the effect of LO on mechanical properties and phase transformation is explored.
In what follows, we refer to calculations using standard IPs with the actual atomic species as TS to differentiate with AA and LOAA simulations that employ an effective potential without species identities.
@@
We begin our study of LOAA with a 2D binary alloy problem. We first consider a random alloy to determine the necessary system size for determining RDFs. Next, we study the effect of LJ parameters on the ordering seen within this alloy. We compare the performance of the standard AA method with the LOAA method to examine the ability of LOAA to capture LO effects. The potential energy per atom and elastic constants of the 2D binary alloy are computed and compared for the TS, AA, and LOAA methods.
III.1.1 2D Lennard-Jones Hexagonal Crystal
@@
$V_{XY}(r)=4\epsilon_{XY}\left[\left(\frac{\sigma_{XY}}{r}\right)^{12}-\left(\frac{\sigma_{XY}}{r}\right)^{6}\right],$ (33)
where $\sigma_{XY}$ and $\epsilon_{XY}$ are LJ parameters for interaction between species $X$ and $Y$ with units of length and energy, respectively, and $X,Y\in\{A,B\}$. For simplicity, we set all $\sigma$ parameters to be the same, $\sigma_{AA}=\sigma_{BB}=\sigma_{AB}=\sigma$, and the potential is truncated at a cutoff radius $r_{\rm cut}$ for all interactions ($AA$, BBBB, ABAB) and shifted to have zero energy at the cutoff, so we have
@@
In simulations, we take $\sigma=1$ and $r_{\rm cut}=2.0$.
We study the effect of a range of $\epsilon_{XY}$ values on the LO. In all cases the ground state crystal structure is hexagonal with a nearest-neighbor distance $a$ that depends on the choice of the $\epsilon$ values. For a crystal containing a single species, the result is $a=1.1159\sigma$.
We begin by examining the radial distribution function for the pure A system. Adapting Eq. (48) in Appendix A to a discrete 2D system, the RDF is
$g(r)=\frac{\Delta n(r)}{2\pi r\Delta r(N/\Omega)},$ (35)
where $N/\Omega$ is the number density (number of atoms $N$ per periodic area $\Omega$), $\Delta r=r_{\rm max}/N_{\rm bins}$, $r_{\rm max}$ is the maximum radius computed, $N_{\rm bins}$ is the number of bins in the histogram used in the computation, and $\Delta n(r)$ is the average number of atoms within a distance $r,r+\Delta r]$ of any other atom,
$\Delta n(r)=\frac{1}{N}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\mathcal{H}((r_{ij}-r)(r+\Delta r-r_{ij})),$ (36)
where $\mathcal{H}(\cdot)$ is the Heaviside step function. It is important to note that the results of the calculation in Eq. (35) depend on the choice of $\Delta r$ (the smaller $\Delta r$, the larger $g(r)$).
In the results below, $r_{\rm max}=r_{\rm cut}$ and $N_{\rm bins}=100$, so that $\Delta r=2.0/100=0.02$. For a hexagonal crystal, $N/\Omega=(3\times 1/6)/(\sqrt{3}a^{2}/4)=2/(\sqrt{3}a^{2})$, which for $a=1.1159\sigma$ (where $\sigma=1$) is $N/\Omega=0.9273$. The RDF for the pure $A$ crystal is show in Fig. 1. There are three peaks in the range $0,r_{\rm cut}]$ associated with the first, second and third neighbor shells at $r=a=1.1159$, $r=\sqrt{3}a=1.9328$ and $r=2a=2.2318$, which each have 6 neighbors.
Figure 1: RDF for a pure AA hexagonal crystal structure.
Next we consider a binary alloy with concentrations $c_{A}=0.8$ and $c_{B}=0.2$. We begin with the random alloy case where the species of any atom is independent of its environment and is set randomly based on the probabilities defined by the concentration values. Since there are now multiple species, we compute partial RDFs,
$g_{XY}(r)=\frac{\Delta n_{XY}(r)}{2\pi r\Delta r(N_{Y}/\Omega)},$ (37)
where $N_{Y}$ is the number of atoms of species $Y$, and $\Delta n_{XY}(r)$ is the average number of atoms of species $Y$ within a distance $r,r+\Delta r]$ of an atom of species $X$,
$\Delta n_{XY}(r)=\frac{1}{N_{X}}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}s_{i}^{X}s_{j}^{Y}\mathcal{H}((r_{ij}-r)(r+\Delta r-r_{ij})).$ (38)
Figure 2: Normalized partial RDF versus number of atoms in the system at the (a) first, (b) second and (c) third peaks. Calculations performed for system sizes of $4\times 2$ ($N=16$), $16\times 9$ ($N=288$), $32\times 18$ ($N=1152$), $48\times 28$ ($N=2688$), and $64\times 37$ unit cells ($N=4736$), with 2 atoms per cell.
As shown in Appendix A, for the case where the species of the atoms are set independently, we have $g_{XY}(r)=g(r)$. This requirement is helpful in determining a minimal system size to ensure that a single snapshot provides good statistics as assumed in Eq. (14), which leads to the effective potential definition in Eq. (16) that we will use below. Fig. 2 presents $g_{XY}(r)/g(r)$ at the three peaks for different system sizes. We see that the results are largely converged at $N=2688$ atoms, although the $BB$ interactions remain noisier due to the lower concentration of $B$ atoms.
We further perform an energy convergence study to determine the necessary minimum system size as shown in Fig. 3. Increasing from $N=16$ to $N=4736$ atoms, the potential energy per atom is computed for 10 realizations (different random seeds) and standard deviation error bars are plotted in Fig. 2. We again see that the results converge around $N=2688$ atoms. To balance simulation costs with accuracy, the following TS calculations are performed with this system size (2688 atoms).
Figure 3: Potential energy per atom versus system size. Each point is an average over 10 computations with standard deviation error bars. Computed at $(\epsilon_{AB},\epsilon_{BB})=(1.0,0.8)$.
@@ III.1.2 Local Ordering in a 2D LJ Binary Alloy
The propensity to LO depends on the nature of bonding in the material, which in our system is determined by the values of the LJ parameters. The energy of a particular arrangement of $B$ atoms substituted into an $A$ lattice, relative to the energy a pure $A$ lattice, for second-neighbor LJ interactions has the following general form (see Appendix B for the derivation):
$\frac{\Delta E(a_{0};\mathcal{S})}{\epsilon_{AA}}=4C(\mathcal{S})\left[\left(\frac{\sigma}{a_{0}}\right)^{12}-\left(\frac{\sigma}{a_{0}}\right)^{6}\right]+4D(\mathcal{S})\left[\left(\frac{\sigma}{\sqrt{3}a_{0}}\right)^{12}-\left(\frac{\sigma}{\sqrt{3}a_{0}}\right)^{6}\right],$ (39)
where $\mathcal{S}$ is a species arrangement forming an LO pattern (defined by $N_{B}$, the number of $B$ atoms, and the relative positions of the $B$ atoms on the AA lattice), $C(\mathcal{S})$ and $D(\mathcal{S})$ are constants that depend on the number of first- and second-neighbor bonds, respectively, for $AA$, ABAB, and $BB$ interactions, and on the LJ parameter ratios $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$. The LJ length parameters are all set to $\sigma=1$, so that $a_{0}=1.1159$ is the equilibrium lattice constant of a pure $A$ lattice. [^10] Any small relaxations in the vicinity of $B$ atoms that might occur are neglected.
@@
To explore how the LJ parameters affect LO, we perform a systematic analysis of Eq. (39) identifying the ground state LO patterns across a range of LJ $\epsilon$ parameters. We follow the procedure in Algorithm III.1.2. For each value $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$, LO patterns are sought for $N_{B}=2$ to 7 $B$ atoms. [^11] The LO patterns are constructed on a template of a hexagonal lattice centered on (0,0) with all points up to the third-neighbor distance from the origin, i.e. 18 neighboring sites plus the origin site for a total of 19. The upper limit of $N_{B}=7$ is the maximum number of $B$ atoms that can separate on the template beyond the second-neighbor distance (i.e. separate beyond interaction), which allows all possible configurations to be fully explored. For each value of $N_{B}$, all distinct LO pattern permutations (accounting for hexagonal symmetry) up to third-neighbors are generated. For example, for $N_{B}=2$ there are three distinct LO patterns where the atoms are separated by the first-neighbor distance ($a_{0}$), second-neighbor distance ($\sqrt{3}a_{0}$), and third-neighbor distance ($2a_{0}$); for $N_{B}=3$ there are 9; and so on up to 70 configurations for $N_{B}=7$ (see Fig. 4). For each LO pattern $\mathcal{S}$, the coefficients $C(\mathcal{S})$ and $D(\mathcal{S})$ in Eq. (39) are computed using Eq. (65) based on the number of $AA$, ABAB, and $BB$ bonds in the configurations and specified values of $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$.
@@
Fig. 5 (left) shows the resulting “phase diagram” where each point is colored according to the LO pattern corresponding to the ground state. Interestingly, there are only two distinct LO states in the ground state of this 2D binary system. Each point colored green favors $BB$ interactions leading to phase separation in the ground state in which clumps of $B$ atoms form inside the AA lattice (see Fig. 5 (middle)). In the orange region, $BB$ interactions are unfavorable, and a solid solution is formed with $B$ atoms uniformly distributed in the AA lattice with spacing between them that exceeds their range of interaction (see Fig. 5 (middle)). We will refer to these two regions as the phase separation (PS) and solid solution (SS) regions. Along the black line separating the two regions, the energy of all LO patterns are the same and therefore the $B$ atoms arrange randomly in the ground state — this is the ideal random alloy case. This line where no local-ordering occurs (i.e. the ‘no-LO line’) is defined by the equation:
III.1.3 Local Ordering Effect on Ground State RDF and Energy

### v4

@@ I Introduction
The average-atom (AA) approach to modeling random alloys is based on an effective medium approximation in which the species of atoms are taken to be independent and consistent with their concentrations in a crystalline alloy. In this method, each atom is identical and typeless, with an effective atomic interatomic potential (IP) in which interactions between atoms are computed as an average over interactions that account for species identities weighted by their probability of occurrence assuming an ideal random alloy. The AA approach was originally introduced by Smith and Was [smith:was:1989], Ackland and Vitek [ackland:vitek:2009], and Najafabadi et al. [najafabadi:wang:1991], and revived by Curtin and coworkers [varvenne:luque:2016, nohring:curtin:2016]. The method has primarily been applied to alloy systems to study the effect of species concentrations on material properties [jian:xie:2020, zhao:xiong:2021, cheng:yuan:2021, seoane:farkas:2022, liang:goodelman:2023, fang:li:2024]. A recent application area for AA is the class of high-entropy alloys (HEAs) [George2019-pl] or multi-principal element alloys [Yeh2004-dy, Shittu2024-wb]. These are alloys that typically contain five or more species of approximately equal concentrations and which are of interest due to their high strength and high fracture toughness relative to traditional alloys [yeh:chen:2004, gludovatz:hohenwarter:2014, zhang:liu:2024]. Such systems require large simulation cells in order to correctly capture species fractional occupancy and hence benefit from AA, which can provide converged results with much smaller system sizes.
@@
Here we extend the AA formalism to explicitly account for LO effects by including information from partial radial distribution functions (RDFs) in the definition of the effective IP. The new local-order average-atom (LOAA) method is as efficient as AA, but correctly captures the effect of LO on alloy properties. In LOAA, the effective interaction between atoms is weighted based on the probabilities of observing pairs of species at a given distance as described by the partial RDFs, as opposed to ideal random alloy probabilities. We validate the approach by studying the elastic properties of a two-dimensional (2D) binary hexagonal crystal with Lennard-Jones (LJ) interactions, a three-dimensional (3D) Fe(1-x)/2Ni(1-x)/2Crx alloy with embedded-atom method (EAM) interactions, and martensitic phase transformations in off-stoichiometric Ni0.67Al0.33 with EAM interactions. We compare the accuracy of LOAA with AA, and with “true species” (TS) computations, calculations using standard IPs with the actual atomic species identities, averaged over many realizations of the system. We find that LOAA provides a more accurate estimate for the TS case than AA with no increase in computational cost except for the initial one-time generation of the partial RDFs. We observe that while the AA method can only accurately predict properties such as lattice constants, energies, and elastic constants for an ideal random alloy, the LOAA formulation correctly predicts these properties with increasing LO.
@@ II.2 Pairwise Interactions
in which $V_{XY}(r)$ is the pair potential function for species $X$ and $Y$, and $r_{ij}$ is the distance between atoms $i$ and $j$.
First, we consider the random alloy case where atomic species are randomly assigned to atoms consistent with a specific set of concentrations. The expectation of the energy follows by applying the ensemble average in Eq. (3):
@@
In Eq. (7), we recover the AA formalism in which the expectation of the energy of the system is expressed in terms of an effective pair potential between typeless atoms (i.e. with no specified species):
@@
In general, Eq. (9) has to be evaluated explicitly, making this approach somewhat more expensive computationally compared to the underlying pair potentials. [^3] For a system containing $n$ species, there are $n(n+1)/2$ function evaluations (accounting for symmetry). Pair potentials with a finite cut-off scale linearly with the number of atoms, $N$. Thus the AA formulation scales as $O(n^{2}N)$. However, as it is more general (and more efficient) to store pair potentials on a discrete grid that is interpolated on-the-fly at runtime, it is usually possible to pre-compute $V_{\rm eff}$ and therefore recover the same speed and scaling of the original pair potential model. The key benefit of the AA approach is that better statistics are obtained than in a calculation of a single realization of a finite-sized random alloy. One is effectively calculating the average behavior over all such possible realizations.
@@
    For a system with LO, i.e. short-range order, $\lim_{r\to\infty}\Phi_{XY}(r;\mathcal{C})=\delta(r;\mathcal{C})$, where $\delta(r;\mathcal{C})=1$ for all distances present in $\mathcal{C}$ and zero otherwise. This indicates that for atoms in $\mathcal{C}$ that are far apart, the probability that they are of species $X$ and $Y$ reverts to the independent case. [^5] We expect this limit to decay quickly with a dependence on the cutoff radius $r_{\rm cut}^{XY}$ of the pair potential $V_{XY}(r)$.
@@
where $g_{XY}(r_{ij})$ is the partial RDF between species $X$ and $Y$. In general, the RDF is defined in terms of a phase average evaluated over the relevant ensemble, such as the canonical ensemble [tuckerman:2010]. This phase average is evaluated for one specific species arrangement satisfying the concentration constraint in Eq. (1). It can be obtained from a molecular simulation or experiment for a given system at a given temperature $T$. Thus the left-hand side of Eq. (14) is the probability over all species arrangements for a single geometric configuration $\mathcal{C}$, whereas the right-hand side involves all geometric configurations with a single species arrangement $\mathcal{S}$. [^6] This means that each atom is assigned a species, such that Eq. (1) is satisfied, and then the phase average is over all possible arrangements of these atoms. The requirement that this equality holds is in the nature of an ergodic hypothesis. Since $\mathcal{C}$ is a system of infinite extent (see Section II.1), it is assumed that the probability over all species arrangements for a single geometric configuration snapshot at finite temperature (left-hand side of Eq. (14)) is equal to the probability over all geometric configurations for a specified set of species (right-hand side of Eq. (14)).
@@
    Computationally, LOAA is not significantly slower than the AA approach in Eq. (9) since $G_{XY}(R)$ can be precomputed. [^8] In the following, we will consider cases where the reference configuration is a perfect crystal. In those cases, LOAA can be made virtually as efficient as the underlying pair potential by precomputing a discrete set of $V_{\rm eff}$ functions, one for each of the discrete pair distances in the lattice up to a conservative neighbor distance allowing for atomic motion. Then during a run, the computation only requires a pair to be mapped to its reference configuration, the value of $R$ computed, and the appropriate $V_{\rm eff}$ function accessed. We note that if the only objective is to determine ground state and the partial RDFs are not known from experiments, then LOAA does not provide a computational benefit over brute force TS simulations since the RDF needs to be obtained. However once obtained, the effective LOAA potential can be used to perform a variety of computations on smaller systems efficiently.
@@ II.3 EAM-style Embedding Function
The passage for the second to third line follows since the terms $U^{X}(\rho_{i})$ and $s_{i}^{X}$ are independent. This is because the density $\rho_{i}$, given in Eq. (20), depends only on the species of the atoms surrounding atom $i$ and not on its species, so that $U^{X}(\rho_{i})$ is independent of the $s^{X}_{i}$ term that follows it in Eq. (21).
@@
In the event that the second-order term is small compared with the leading term in Eq. (24) (as for example shown for Cu-Ni alloys in [najafabadi:wang:1991]) and ${\rm Var}_{\mathcal{S}}(\rho_{i})$ is small (indicating convergence of the expansion in Eq. (22)), then
@@ III Results
To explore the ability of the LOAA methodology to capture LO effects, we study two example problems: 1) a 2D binary alloy modeled using an LJ potential where the degree of LO can be tuned by the choice of LJ parameters to allow for systematic comparison with the AA approach; and 2) two 3D metallic alloy systems (FeNiCr and NiAl) modeled using EAM potentials where the effect of LO on mechanical properties and phase transformation is explored.
In what follows, we refer to calculations using standard IPs with the actual atomic species as the “true species” (TS) case to differentiate with AA and LOAA simulations that employ an effective potential without species identities.
@@
As a first test case for LOAA, we consider a 2D binary alloy modeled using an LJ potential. We study the effect of the LJ parameters on the short-range order seen within this alloy. The potential energy per atom and elastic constants of the 2D binary alloy are computed and compared for the TS, AA, and LOAA methods, examining the importance of LO effects.
III.1.1 Problem Definition
@@
$V_{XY}(r)=4\epsilon_{XY}\left[\left(\frac{\sigma_{XY}}{r}\right)^{12}-\left(\frac{\sigma_{XY}}{r}\right)^{6}\right].$ (33)
Here $\sigma_{XY}$ and $\epsilon_{XY}$ are LJ parameters for interaction between species $X$ and $Y$ with units of length and energy, respectively, and $X,Y\in\{A,B\}$. For simplicity, we set all $\sigma$ parameters to be the same, $\sigma_{AA}=\sigma_{BB}=\sigma_{AB}=\sigma$, and the potential is truncated at a cutoff radius $r_{\rm cut}$ for all interactions ($AA$, BBBB, ABAB) and shifted to have zero energy at the cutoff. Thus we have
@@
In simulations, we take $\sigma=1$ and $r_{\rm cut}=2.0$. The ground state crystal structure is 2D hexagonal (triangular) with a nearest-neighbor distance $a_{0}$ that depends on the choice of the $\epsilon_{XY}$ values. For a crystal containing a single species, the result is $a_{0}=1.1159\sigma$.
@@ III.1.2 Local Ordering in a 2D LJ Binary Alloy
In general, the propensity to LO depends on the nature of bonding in the material. In the present case this is determined by the values of the LJ parameters defined in Section III.1.1. An analysis of the energetics of 2D binary hexagonal crystals for second-neighbor LJ interactions (given in Appendix B) shows that the energy of a particular arrangement of $B$ atoms substituted into an $A$ lattice relative to the energy of a pure $A$ lattice has the following general form:
$\frac{\Delta E(a_{0};\mathcal{S})}{\epsilon_{AA}}=4C(\mathcal{S})\left[\left(\frac{\sigma}{a_{0}}\right)^{12}-\left(\frac{\sigma}{a_{0}}\right)^{6}\right]+4D(\mathcal{S})\left[\left(\frac{\sigma}{\sqrt{3}a_{0}}\right)^{12}-\left(\frac{\sigma}{\sqrt{3}a_{0}}\right)^{6}\right].$ (35)
Here $\mathcal{S}$ is a species arrangement forming an LO pattern (defined by the number of $B$ atoms $N_{B}$, and the relative positions of the $B$ atoms on the AA lattice), $C(\mathcal{S})$ and $D(\mathcal{S})$ are constants that depend on the number of first- and second-neighbor bonds, respectively, for $AA$, ABAB, and $BB$ interactions, and on the LJ parameter ratios $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$, and $a_{0}=1.1159$ as noted above. [^9] Any small relaxations in the vicinity of $B$ atoms that might occur are neglected.
@@
To explore how the LJ parameters affect LO, we perform a systematic analysis of Eq. (35) identifying the ground state LO patterns across a range of LJ $\epsilon$ parameters. We follow the procedure in Algorithm III.1.2. For each value $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$, LO patterns are sought for $N_{B}=2$ to 7 $B$ atoms. [^10] The LO patterns are constructed on a template of a hexagonal lattice centered on (0,0) with all points up to the third-neighbor distance from the origin, i.e. 18 neighboring sites plus the origin site for a total of 19. The upper limit of $N_{B}=7$ is the maximum number of $B$ atoms that can separate on the template beyond the second-neighbor distance (i.e. separate beyond interaction), which allows all possible configurations to be fully explored. For each value of $N_{B}$, all distinct LO pattern permutations (accounting for hexagonal symmetry) up to third-neighbors are generated. [^11] For a second-neighbor cutoff, it is necessary to consider atoms at third neighbor distance to include patterns in which one or more atoms are beyond the range of interaction. For $N_{B}=2$ there are three distinct LO patterns where the atoms are separated by the first-neighbor distance ($a_{0}$), second-neighbor distance ($\sqrt{3}a_{0}$), and third-neighbor distance ($2a_{0}$), for $N_{B}=3$ there are 9, and so on up to 70 configurations for $N_{B}=7$ (see Fig. 1). For each LO pattern $\mathcal{S}$, the coefficients $C(\mathcal{S})$ and $D(\mathcal{S})$ in Eq. (35) are computed using Eq. (56) based on the number of $AA$, ABAB, and $BB$ bonds in the configurations and specified values of $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$.
@@
Fig. 2 (left) shows the resulting “phase diagram” where each point is colored according to the LO pattern corresponding to the ground state. Interestingly, there are only two distinct LO states in the ground state of this 2D binary system. Each point colored green favors $BB$ interactions leading to phase separation in the ground state in which $B$ atoms group together within the AA lattice (see Fig. 2 (middle)). In the orange region, $BB$ interactions are unfavorable, and a solid solution is formed with $B$ atoms separated by a distance exceeding their range of interaction (see Fig. 2 (middle)). We will refer to these two regions as the phase separation (PS) and solid solution (SS) regions. Along the black line separating the two regions, the energy of all LO patterns are the same and therefore the $B$ atoms arrange randomly in the ground state — this is the ideal random alloy case. We refer to this as the ‘no-LO line’ — see Appendix C for a more detailed discussion of the nature of this line.
III.1.3 Monte Carlo Simulations of the Ground State